# On Creativity and Open-Endedness

### v1

@@ Abstract
ALife (ALife) as an interdisciplinary field draws inspiration and influence from a variety of perspectives. Scientific progress crucially depends, then, on concerted efforts to invite cross-disciplinary dialogue. The goal of this paper is to revitalize discussions of potential connections between the fields of CC (CC) and ALife, focusing specifically on the concept of OE (OE); the primary goal of CC is to endow artificial systems with creativity, and ALife has dedicated much research effort into studying and synthesizing OE and artificial innovation. However, despite the close proximity of these concepts, their use so far remains confined to their respective communities, and their relationship is largely unclear. We provide historical context for research in both domains, and review the limited work connecting research on creativity and OE explicitly. We then highlight specific questions to be considered, with the eventual goals of (i) decreasing conceptual ambiguity by highlighting similarities and differences between the concepts of OE, (ii) identifying synergy effects of a research agenda that encompasses both OE and creativity, and (iii) establishing a dialogue between ALife and CC research.
@@ Introduction
The engineering of both CC (CC) and OE (OE) have equally been advocated as the “final frontier” [Colton et al., 2012], and the “last grand challenge” [Stanley et al., 2017] of AI (AI) research, respectively. These claims, together with an immense and growing body of related work, highlight the magnitude of, and desire toward accomplishing, these two research challenges. They constitute central and enduring research themes found in creativity, ALife, and AI that are addressed, amongst others, by the CC and ALife (ALife) communities independently. We note that neither challenge is close to being fully solved, but both overlap in their overarching purpose. We consequently ask: Are CC and OE entirely different concepts, and does their synthesis represent disconnected research endeavors? Are the concepts within both mutually exclusive or do they overlap (and to what extent)?
@@
There is much reason to believe that this is not the case. Creativity, more broadly, is understood as a hallmark feature of human and (with nuances) animal cognition, while OE is considered foundational for the evolution and development of advanced lifeforms. OE is often used to characterize tasks and tests that assess human creativity, and as a challenge to motivate the advancement of CC algorithms. Both creativity and OE have been proposed together as features of (especially evolutionary) algorithms across both disciplines.
@@
We argue that these similarities signal opportunity: there is potential for the CC and ALife communities to learn about each other’s research heritage and challenges, and, especially in close interaction, use it to advance research on either front. Making use of this opportunity is important, since research involving both creativity and OE has unresolved ambiguities in the involved concepts, including their definitions and measurements. In a way, this also marks an (unfortunate) commonality between the two areas.
@@
The premise of our paper is threefold: (1) to ask questions on specific connections between CC and OE that have the potential to promote further work, (2) to identify synergy effects of a research agenda that encompasses both, and (3) to consequently spark discussion and collaboration between the CC and ALife communities. To this end, we bring together a team of CC and ALife researchers to make two distinct contributions: we (a) summarise connections from isolated, existing work on both concepts and (b) uncover new connections based on our expert discussions. While our goal is ultimately to deepen our understanding of each concept and the requirements to reproduce them in artificial systems, we do not pretend to resolve these connections conclusively within this limited space and team of authors. Instead, we ask provoking questions on the nature and constituents of these connections to establish a starting point for additional deep, dedicated interdisciplinary investigations.
@@
By doing so, we continue and revitalize the discussion proposed in Soros and Stanley [2016], asking “Are there nontrivial differences between creativity and OEE, or are they fundamentally the same concept?” This questions whether creativity and OEE have nontrivial differences or are fundamentally the same concept. Existing work in CC focuses on OE as a property of tasks / exercises / problems [e.g. Epstein, 2011] or environments [e.g. Guckelsberger et al., 2016], but not as a characteristic of the process or the resulting product which ALife seeks to reproduce. Vice versa, ALife researchers have drawn connections from OE to creativity, but articulated only as sidelines. Instances of work spanning both fields and research agendas [Boden, 2009; Soros and Stanley, 2016] are promising, but few and isolated.
@@
While this paper was written primarily with the goals and understanding of CC and ALife researchers in mind, creativity and OE are becoming dramatically more relevant beyond the boundaries of these communities. We hope that our insights in this paper can inspire and advance AI more broadly.
@@
The concepts of OE and creativity (more broadly) are used in a multitude of fields and contexts, such as educational research & psychology, games & play, software engineering, and design, just to name a few. Here, we focus on their use specifically within the fields of ALife (ALife) and CC (CC), separate from adjacent fields such as machine learning and AI research more generally. At the same time, we acknowledge that the boundaries between fields are fluid and evolving [^1] Cf. e.g. the workshops on “Agent Learning in Open-Endedness” (ALOE) and “Intrinsically Motivated Open-ended Learning” (IMOL) at NeurIPS, which also feature ALife research.. As a first notable difference between both concepts, they play a very different role in the definition of the respective discipline’s identity.
@@
CC has been defined as “the art, science, philosophy, and engineering of computational systems which, by taking on particular responsibilities, exhibit behaviors that unbiased observers would deem to be creative” [Colton et al., 2012]. This places creativity right at the core of the discipline’s identity. Researchers follow two broad goals: “understanding human creativity” and “producing machine creativity” [Boden, 2004; Pérez y Pérez, 2018]. The latter can be differentiated into co-creative systems [Kantosalo and Toivonen, 2016], taking on some creative responsibility [Colton et al., 2012] in interaction with others, and fully autonomous creative systems [Colton, 2008].
CC’s conceptualization of creativity is mostly adopted from (psychological) creativity studies, the long-standing joint effort of psychology, sociology, philosophy, and other non-computational disciplines to investigate creativity. CC’s ontological contribution then consists of re-evaluating and extending the scope and attribution of creativity in the context of (typically computational) artificial systems. Due to its adoption from creativity studies, CC follows a strongly anthropocentric agenda. As a result, creativity is almost exclusively understood in terms of human creativity.
@@
Runco and Jaeger [2012] have helped mitigate this situation by abstracting common underlying factors in existing studies and definitions, arriving at a two-component “standard definition” of creativity which requires a product or process to be both novel and valuable to be deemed creative. Novelty can be minimal and is often not considered sufficient; the definition is extended by a third component– surprise [Boden, 2004]. With the surge of more powerful creative AI systems, this “standard definition” has recently been updated to fence human from artificial creativity via the features of authenticity and intentionality [Runco, 2023]. Crucially though, this proposal can be argued to lack behind CC research which has long embraced both features as part of its central challenge [cf. Ventura, 2016; Guckelsberger et al., 2017; Colton et al., 2018; McCormack et al., 2019, e.g.].
Unfortunately, these definitions only resolve some of the ambiguity while introducing more. While the meaning of authenticity and intentionality remains subject of philosophical debate, the meanings of value, novelty, and surprise vary depending on who evaluates the creativity, of what, and in which context [Silvia, 2018; Puryear and Lamb, 2020]. As a consequence of this conceptual ambiguity, domain-, observer- and perspective-sensitivity, no consensus exists on standardised measures to assess CC [Lamb et al., 2018], despite various efforts [e.g. Ritchie, 2019].
Artificial Life, Open-Ended Evolution, and Eventual Open-Endedness
@@
It is important to note that the concept of open-ended evolution (OEE), arising from the ALife community predates the generalization of broad open-endedness (OE). Foundational descriptions of OEE (OEE) originate from many pre-millennium works. Specifically, Bedau [1998] finds the phrase “open-ended evolution” previously used by Lindgren [1991] and Ray [1996]. Bedau et al. [2000] then establishes determining “what is inevitable in the open-ended evolution of life” as a grand challenge for ALife. After nearly two decades of work on the problem, Packard et al. [2019] note that evolution’s dynamic production of diversity “has led life’s creative productivity to be called Open-Ended Evolution (OEE) in the field”. The typical approach towards understanding and synthesizing the phenomenon has been to create virtual worlds imbued with features of the natural world thought to be key to unbounded generation (of whatever things, be it code snippets or more physically embodied agents), and then performing systemic experiments probing exactly how long the generative process lasts and for what reasons it might fail. Notable systems establishing the foundation for current research include Tierra [Ray, 1992], Avida [Ofria and Wilke, 2004], PolyWorld [Yaeger, 1994], Geb [Channon and Damper, 2000], Cosmos Taylor [2004], Division Blocks [Spector et al., 2007], Evosphere [Miconi and Channon, 2005], and Chromaria [Soros and Stanley, 2014]. It should be noted that this list is not exhaustive.
Throughout these decades, many definitions of OEE have been proposed. To only name a few, it has been defined as:
  • a system in which the “number of possible types by far exceeds the number of individuals (copies, sequences, etc.) in a plausible (realistic) population” [Maynard-Smith and Szathmary, 1995],
  • the “on-going and indefinitely creative production of significantly new kinds of adaptive responses to significantly new kinds of adaptive challenges and opportunities” [Bedau, 1998],
  • “a system in which components continue to evolve new forms continuously, rather than grinding to a halt when some sort of ‘optimal’ or stable position is reached” [Taylor, 1999], as well as
@@
Dolson et al. [2015] turn the question of what OEE is (and, accordingly, how it should be measured) on its head and ask instead what OEE is not. Key indicators include a lack of change in the population, a lack of novelty, a lack of increasing complexity, a lack of ecosystem diversity, and a lack of major transitions. This account then admits a pluralistic framing of the phenomenon, allowing us to consider a wide variety of systems ranging from definitely-not-open-ended to very-strongly-open-ended. However, the steps in this gradient (and the gradient in general) remain unclear and difficult to categorize at present.
@@
The idea of OE as a property not exclusive to evolution has been a topic of exploration in the years bridging the gap between the field’s inception and its more recent diffusion into adjacent fields. Lehman et al. [2008] establish “a link between the challenge of solving highly ambitious problems in machine learning and the goal of reproducing the dynamics of open-ended evolution in artificial life”. They refer to “open-endedness in natural evolution”, hinting at the possibility that other kinds of open-endedness might exist apart from the biological or Darwinian kind. Pattee and Sayama [2019] note that “open-endedness is often considered a prerequisite property of the whole evolutionary system and its dynamical behaviors”, explicitly separating the concepts of OE and evolution.
Much like CC, the idea of OE is enveloped in ambiguous concepts (like novelty and unboundedness, for example) that are difficult to measure quantitatively and are nearly impossible to untangle from context. Even so, researchers in ALife seem to agree on a broad definition of OE, even if the details vary. In general, OE is defined as the ability to continually create new, innovative structures and evolve under new, innovative dynamics, all while continually increasing in complexity [Banzhaf et al., 2016].
@@
Similar to the case of creativity, beyond conceptual clarity, there is a distinct lack of consensus for “gold-standard” quantitative measures for open-endedness [Taylor et al., 2016]. This makes it difficult to interpret the resulting behaviors/dynamics of new, innovative ALife models that agree with a field-wide consensus. Because of this, there is a strong push from within the ALife community to further explore these concepts [Song, 2022]. Without this consensus, it remains an open challenge to 1) develop a theory for open-endedness that easily translates to measures and algorithmic design in silico and 2) imitate and recreate real open-ended systems like biological evolution artificially.
@@
As our first contribution, we provide an up-to-date review of existing connections between OE and creativity drawn in related work. First though, we highlight which work has been excluded from this review.
@@
Explicit discussions on the relationship of creativity and OE are scarce and we exclude any work that merely draws implicit connections from this review. For instance, Stanley et al. [2017] do not directly compare creativity and OE, but on two occasions amalgamate the concepts into “open-ended creativity ”. Open-endedness is framed as a modality of a process (creativity) that could just as well be bounded, akin to traditional problem-solving. Similarly, [2020] asserts that evolution is inherently creative, but lacks the foresight and intentionality of human creativity. Lehman et al. [2020] and Wiggins et al. [2015] draw connections between creativity and evolution, but neither of these investigations connects creativity with OE as a property of evolutionary processes. As a final example, Boden [2015], who has had a major influence on the development of CC as a research endeavor, finds parallels between three types of creative processes (combinatory, exploratory, and transformational) and the research matter of ALife, but does not relate these to OE specifically. This has later been mended by Taylor [2019, 2021], as discussed below.
We also exclude work that focuses on OE as a property of tasks/exercises/problems – which have more than one right answer, solution or outcome and can be completed in different ways [e.g. Epstein, 2011]. Similarly, we exclude work on open-ended environments [e.g. Guckelsberger et al., 2016] within which such tasks can be established and solved. Open-ended tasks are often cited as a necessary condition for the emergence of creativity. Such tasks can be completed in different ways, and are distinct from OE as characteristic of processes addressed in ALife (see Background). This exclusion rests on our understanding that connections between creativity and OE as a subject of CC and ALife research can best be drawn by focusing on the fields’ effort to synthesize a creative or open-ended process. This process binds agent embodiment and environment together in one system and is potentially constrained by a task or giving rise to tasks (cf. creativity as problem-finding). In this view, agent embodiment, environment, and task represent variables to e.g. benchmark the generality of this synthesized process against. We leverage literature on creativity and OE as a process to anchor our investigation and discuss the systems perspective on creativity and OE as a conceptual commonality below. Notably, this excludes literature from educational and creativity research investigating the potentiality of different tasks and environments to assess and exercise human creativity.
@@
Boden [2015], mentioned previously, provides a detailed conceptual connection between CC and ALife. First, “psychological creativity” (involving human thought and activity) is contrasted with “biological creativity”, which is described as “the ability to generate new cells, organs, organisms, or species. This includes the (phylogenetic) creativity of evolution, the (ontogenetic) creativity of morphogenesis, and the (autopoietic) creation of individual cells”. Then, as mentioned previously, creativity in the broader sense is classified into three types: “the different sorts of surprise that we experience on encountering a creative idea correspond to different mechanisms for producing novelty. Those mechanisms mark three kinds of creativity: combinational, exploratory, and transformational.”
Soros and Stanley [2016] additionally drew connections soon thereafter, focusing more specifically on OEE. To this end, they highlight a potential connection between Taylor et al. [2016]’s primary categories of hallmarks of open-endedness (ongoing adaptive novelty, and ongoing growth of complexity) and Boden’s concepts of exploratory and transformational creativity [Boden, 2004]. Additionally, they raise three open questions about the relationship between OEE and creativity theory: 1) Are there nontrivial differences between creativity and OEE or are they fundamentally the same concept? 2) Is there any utility in formally pursuing the nature of the relationship between evolution and creativity? 3) Do there exist metrics and practices for studying computational creativity (or creativity in general) that could help identify and quantify hallmarks of open-ended evolution? They thus appeal to the struggles in both field to define and measure creativity and OE, respectively.
Taylor [2019] discusses three different kinds of novelties described by [Banzhaf et al., 2016], based on three different types of creativity (combinatorial, exploratory, and transformational) proposed by [Boden, 2004, 2015]: exploratory, expansive, and transformational novelty. Combinatorial creativity refers to the creation of new concepts or artifacts by combining features of existing ones. Exploratory creativity corresponds to traversing a space of such concepts or artifacts to construct or discover new and valued instances. Finally, transformational creativity is defined as the reaching of new points that only become accessible by altering the rules defining the space itself “so that thoughts are now possible which previously (…) were literally inconceivable” [Boden, 2004, p. 6].

### v2

@@ Abstract
ALife (ALife) as an interdisciplinary field draws inspiration and influence from a variety of perspectives. Scientific progress crucially depends, then, on concerted efforts to invite cross-disciplinary dialogue. The goal of this paper is to revitalize discussions of potential connections between the fields of CC (CC) and ALife, focusing specifically on the concept of OE (OE); the primary goal of CC is to endow artificial systems with creativity, and ALife has dedicated much research effort into studying and synthesizing OE and artificial innovation. However, despite the close proximity of these concepts, their use so far remains confined to their respective communities, and their relationship is largely unclear. We provide historical context for research in both domains, and review the limited work connecting research on creativity and OE explicitly. We then highlight specific questions to be investigated in future work, with the eventual goals of (i) decreasing conceptual ambiguity by highlighting similarities and differences between the concepts of OE, (ii) identifying synergy effects of a research agenda that encompasses both OE and creativity, and (iii) establishing a dialogue between ALife and CC research.
@@ Introduction
This paper seeks connections between two fields, communities, and research endeavours: ALife (ALife) and CC (CC). ALife researchers study “life and life-like processes through simulation and synthesis” [Bedau, 2007] in software, hardware, and wetware, whereas CC researchers are concerned with the theory and engineering of computational systems capable of exhibiting creativity, including systems that co-create with others [e.g. Colton et al., 2012]. With regards to ALife, we focus on what has been considered a grand challenge [Bedau et al., 2000] and major goal [Stepney, 2021] of the field: the definition, categorization, and measurement of OE (OE) as a characteristic of life and natural evolution.
@@
The engineering of both CC and OE have equally been advocated as the “final frontier” [Colton et al., 2012], and the “last grand challenge” [Stanley et al., 2017] of AI (AI) research, respectively. In AI research more generally, OE has been claimed to be essential for any artificial superhuman intelligence [Hughes et al., 2024]. These claims, together with an immense heritage and growing body of related work, highlight the magnitude of, and desire toward accomplishing, these two research challenges. At the basis of this paper lies the observation that neither challenge is close to being fully solved, but, at the same time, both challenges appear related. We consequently ask: Are creativity and OE entirely different concepts, or do their synthesis and measurement represent closely related research endeavors? And if so, how are they connected?
@@
There is much reason to believe in the latter. Creativity, more broadly, is understood as a hallmark feature of human [Still and d’Inverno, 2016] and (with nuances) animal cognition [Kaufman and Kaufman, 2015], while OE is considered foundational for the evolution and development of advanced lifeforms [Bedau et al., 2000]. In creativity studies [Williams et al., 2016], OE is often used as a property of tasks and tests to assess human creativity [Epstein, 2011] and in CC, it is used as challenge to motivate the advancement of algorithms [e.g. Guckelsberger et al., 2016]. Both creativity and OE have been proposed together as features of (especially evolutionary) algorithms across both disciplines [e.g. Stanley and Lehman, 2015]. Moreover, creativity is very often mentioned together with, or amalgamated in, definitions of OE [Bedau, 1998; Packard et al., 2019; Hughes et al., 2024], albeit without further differentiation.
@@
We argue that these observations signal opportunity: there is potential for the CC and ALife communities to learn about each other’s research heritage and challenges, and, especially in close interaction, use it to advance research on either front. Making use of this opportunity is important, since research involving both creativity and OE has unresolved ambiguities in the involved concepts, including their definitions and measurements. In a way, this also marks an (unfortunate) commonality between the two areas.
@@
For this paper, we brought together a team of CC and ALife researchers to make three distinct contributions: we (1) summarise previously identified connections between CC and OE from isolated, existing work; we (2) uncover new connections that have the potential to promote further work based on our expert discussions; and we (3) identify synergy effects of a research agenda that encompasses both with the goal to spark discussion and interdisciplinary collaboration between CC, ALife, and other communities.
While our goal is ultimately to deepen our understanding of each concept and the requirements to reproduce them in artificial systems, we do not pretend to resolve these connections conclusively given limited space, time, and expertise. Instead, we ask provoking questions on the nature and constituents of these connections to establish a starting point for additional deep, dedicated interdisciplinary investigations.
@@
By doing so, we continue and revitalize a prior discussion by Soros and Stanley [2016], asking “Are there nontrivial differences between creativity and OEE, or are they fundamentally the same concept?” Our treatment is more extensive and general, as it also considers OE beyond OEE (OEE). Existing work in CC focuses on OE as a property of tasks / exercises / problems [e.g. Epstein, 2011] or environments [e.g. Guckelsberger et al., 2016], but not as a characteristic of the process or the resulting product which ALife seeks to reproduce. Vice versa, ALife researchers have drawn connections from OE to creativity, but articulated only as sidelines. Instances of work spanning both fields and research agendas [Boden, 2009; Soros and Stanley, 2016] are promising, but few and isolated.
@@
While this paper was written primarily with the goals and understanding of CC and ALife researchers in mind, creativity and OE are becoming dramatically more relevant beyond the boundaries of these communities, especially for AI more broadly [e.g. Hughes et al., 2024]. This paper is written for a wider audience, and we hope that our insights can inspire and advance research in AI and beyond.
@@
The concepts of OE and creativity (more broadly) are used in a multitude of fields and contexts, such as educational research & psychology, games & play, software engineering, and design, just to name a few. Here, we focus on their use specifically within the fields of ALife (ALife) and CC (CC), separate from adjacent fields such as machine learning and AI research more generally. At the same time, we acknowledge that the boundaries between fields are fluid and evolving [^1] E.g. the workshops on “Agent Learning in Open-Endedness” (ALOE) and “Intrinsically Motivated Open-ended Learning” (IMOL) at NeurIPS also feature CC and ALife research.. As a first notable difference between both concepts, they play a very different role in the definition of the respective discipline’s identity. We next introduce each discipline in more detail, together with the core concepts of creativity, OE and OEE.
@@
CC has been defined as, amongst others, “the art, science, philosophy, and engineering of computational systems which, by taking on particular responsibilities, exhibit behaviors that unbiased observers would deem to be creative” [Colton et al., 2012]. This places creativity right at the core of the discipline’s identity. Researchers follow two broad goals: “understanding human creativity” and “producing machine creativity” [Boden, 2004; Pérez y Pérez, 2018]. The latter can be differentiated into co-creative systems [Kantosalo and Toivonen, 2016], taking on some creative responsibility [Colton et al., 2012] in interaction with others, and fully autonomous creative systems [Colton, 2008] – marking two extremes on an interaction continuum. A deeper introduction of CC as a field is beyond the scope of this paper, and we refer the reader to excellent overview articles of its early history and academic predecessors [Boden, 2015b] as well as more recent depiction [Veale et al., 2019]. Crucially, the goals introduced above are increasingly also pursued in other fields and under different labels such as Creative AI and Generative AI, amongst others [Vear and Poltronieri, 2022].
CC’s conceptualization of creativity is mostly adopted from (psychological) creativity studies, the long-standing joint effort of psychology, sociology, philosophy, and other non-computational disciplines to investigate creativity [Williams et al., 2016]. CC’s ontological contribution then consists of re-evaluating and extending the scope and attribution of creativity in the context of (typically computational) artificial systems. Similarly to creativity studies, CC follows an anthropocentric agenda, and creativity is almost always connected to features of human creativity.
@@
Runco and Jaeger [2012] have helped mitigate this situation by abstracting common underlying factors in existing studies and definitions, arriving at a two-component “standard definition” of creativity which requires a product or process to be both novel and valuable to be deemed creative. Novelty can be minimal and may not be considered sufficient; the definition is consequently sometimes extended by a third component– surprise [Boden, 2004]. With the surge of more powerful creative AI systems, this “standard definition” has recently been updated to fence human from “artificial creativity” via the features of authenticity and intentionality [Runco, 2023]. This proposal can be argued to lack behind CC research which has long recognised both features as central challenges and identified algorithmic proposals for their realisation [cf. Ventura, 2016; Guckelsberger et al., 2017; Colton et al., 2018; McCormack et al., 2019, e.g.].
Unfortunately, these definitions only resolve some of the ambiguity while introducing more. In particular the meaning of authenticity and intentionality remain subject of philosophical debate, and the meanings of value, novelty, and surprise vary depending on who evaluates the creativity, of what, and in which context [Silvia, 2018; Puryear and Lamb, 2020]. As a consequence of this conceptual ambiguity, domain-, observer- and perspective-sensitivity, no consensus exists on standardised measures to assess CC [Lamb et al., 2018], despite various efforts [e.g. Ritchie, 2019].
ALife, Open-Ended Evolution, and Open-Endedness
@@
It is important to note that the concept of open-ended evolution (OEE), arising from the ALife community, predates the generalization to OE (OE) more broadly. Foundational descriptions of OEE originate from many pre-millennium works. Specifically, Bedau [1998] finds the phrase “open-ended evolution” previously used by Lindgren [1991] and Ray [1996]. Bedau et al. [2000] then establishes determining “what is inevitable in the open-ended evolution of life” as a grand challenge for ALife. Although several more grand challenges have been identified, this still emphasises the importance of OE and OEE to ALife research. Notably different to the concept of creativity in CC though, OE/OEE are better understood as supporting ALife’s central goal of studying and replicating life and life-like processes, rather than as the central goals themselves.
After nearly two decades of work on the problem, Packard et al. [2019] note that evolution’s dynamic production of diversity “has led life’s creative productivity to be called Open-Ended Evolution (OEE) in the field” [^2] The curious reader is directed towards Packard et al. [2019] for an excellent and more extensive introduction to OEE (OEE) and OE (OE).. The typical approach towards understanding and synthesizing the phenomenon has been to create virtual worlds imbued with features of the natural world thought to be key to unbounded generation (of whatever things, be it code snippets or embodied agents), and then performing systemic experiments probing exactly how long the generative process lasts and for what reasons it might fail. Notable systems establishing the foundation for current research include Tierra [Ray, 1992], Avida [Ofria and Wilke, 2004], PolyWorld [Yaeger, 1994], Geb [Channon and Damper, 2000], Cosmos [Taylor, 2004], Division Blocks [Spector et al., 2007], Evosphere [Miconi and Channon, 2005], and Chromaria [Soros and Stanley, 2014]. It should be noted that this list is not exhaustive.
Throughout these decades, many definitions of OEE (but, to the best of our knowledge, no definitions of OE) have been proposed. To only name a few, it has been defined as:
  • a system in which the “number of possible types by far exceeds the number of individuals (copies, sequences, etc.) in a plausible (realistic) population”
    [Maynard-Smith and Szathmáry, 1995],
  • the “on-going and indefinitely creative production of significantly new kinds of adaptive responses to significantly new kinds of adaptive challenges and opportunities”
    [Bedau, 1998],
  • “a system in which components continue to evolve new forms continuously, rather than grinding to a halt when some sort of ‘optimal’ or stable position is reached”
    [Taylor, 1999]
@@
Dolson et al. [2015] turn the question of what OEE is (and, accordingly, how it should be measured) on its head and ask instead what OEE is not. Key indicators include a lack of change in the population, a lack of novelty, a lack of increasing complexity, a lack of ecosystem diversity, and a lack of major transitions. This pluralistic account then admits more differentiation, allowing us to consider a wide variety of systems ranging from definitely-not-open-ended, lacking all of these features, to very-strongly-open-ended. However, the steps in this gradient (and the gradient in general) remain unclear and difficult to categorize at present.
@@
The remainder of this section shifts from discussing OEE specifically to discussing OE in the broader sense. As noted earlier in this section, OE can be considered a property of systems that might be evolutionary or not. A few notable works functionally decouple OE and evolution in their definitions. Lehman et al. [2008] establish “a link between the challenge of solving highly ambitious problems in machine learning and the goal of reproducing the dynamics of open-ended evolution in artificial life”. They refer to “open-endedness in natural evolution”, hinting at the possibility that other kinds of open-endedness might exist apart from the biological or Darwinian kind. At the same time, Miconi [2008], writing about coevolution, refers to the phenomenon as “the simplest way to implement a more natural, possibly open-ended form of evolution”. What is significant about this framing is the implicit statement that not all kinds of evolution are necessarily open-ended. Pattee and Sayama [2019] additionally note that “open-endedness is often considered a prerequisite property of the whole evolutionary system and its dynamical behaviors”, further explicitly separating the concepts of OE and evolution.
@@
Much like creativity, the concept of OE in ALife is ambiguous and, arguably even more so than creativity in CC, its definition is subject to ongoing debate. Banzhaf et al. [2016] summarize historical disagreements on what constitutes OE, noting that definitions tend to focus on the “continual (unbounded) creation” of either novelty or complexity. Also similar to creativity in CC though, these definitional components (e.g. novelty) are in themselves ambiguous and nearly impossible to untangle from context. As another similarity, there is a distinct lack of consensus for “gold-standard” quantitative measures for open-endedness [Taylor et al., 2016]. This makes it difficult to identify and consequently analyze the behaviors/dynamics of new, innovative ALife models that agree with a field-wide consensus. Song [2022] notes the “disjunction [in ALife] between the conceptual definitions, operational definitions, and the actual implementation of measurement” of OE. As such, it remains an open challenge to (1) develop a theory for open-endedness that easily translates to measures and algorithmic design in silico and (2) imitate and recreate real open-ended systems like biological evolution artificially.
Some of the ambiguity in conceptualisation and measurement may not be symmetrical between ALife and CC, offering an opportunity to mitigate these challenges in ALife, and similarly address open questions in CC. To exploit these opportunities though, we must better understand the connection between creativity and OE/OEE. This need is also demonstrated by the undifferentiated mentioning of creativity in writing about [e.g. Packard et al., 2019] and definitions of OE/OEE [e.g. Bedau, 1998], also quoted in this section.
@@
As our first contribution, we provide an up-to-date review of connections between OE and creativity previously drawn in related work. First, though, we highlight the scope of this review and which work has been excluded.
@@
Explicit discussions on the relationship of creativity and OE are scarce and we exclude any work that merely draws implicit connections and does not explicate the involved concepts from this review. For instance, Stanley et al. [2017] do not directly compare creativity and OE, but on two occasions amalgamate the concepts into “open-ended creativity”. OE is framed as a modality of a process (creativity) that could just as well be bounded, akin to traditional problem-solving. Similarly, Lehman et al. [2020] assert that evolution is inherently creative, but lacks the foresight and intentionality of human creativity. Here, OE is not being directly related to creativity, and human creativity is not further defined. Lehman et al. [2020] and Wiggins et al. [2015] draw connections between creativity and evolution, but neither of these investigations connects creativity with OE as a property of evolutionary processes. As a final example, Boden [2015a], who has had a major influence on the development of CC as a research endeavor, finds parallels between three types of creative processes (combinational, exploratory, and transformational – defined later) and the research matter of ALife, but does not relate these to OE specifically. As discussed below, this was later mended by Taylor [2019, 2021].
We also exclude work that focuses on OE as a property of tasks/exercises/problems – which have more than one right answer, solution or outcome and can be completed in different ways [e.g. Epstein, 2011]. Open-ended tasks are often cited as a necessary condition for the emergence of creativity. Closely related, open-ended environments [e.g. Guckelsberger et al., 2016] can be considered as the frames within which such tasks can be established and solved. We exclude both types of related work as they are not compatible with our focus on OE as a process (see below); we consider a task akin to an objective which might be defined a priori by an observer external to the system such as a person aiming to assess the other’s creativity. We argue that connections between creativity and OE as a subject of CC and ALife research can best be drawn by focusing on the fields’ effort to synthesize a creative or open-ended process (see Background). This process binds agent embodiment and environment together in one system and is potentially constrained by a task or giving rise to tasks (cf. creativity as problem-finding). In this view, agent embodiment, environment, and task represent variables to e.g. benchmark the generality of this synthesized process against. We leverage literature on creativity and OE as a process to anchor our investigation and discuss systems perspectives on creativity and OE as a conceptual commonality below. As a consequence, we notably exclude literature from educational studies and creativity research investigating the potentiality of different tasks and environments to assess and exercise human creativity.
@@
Boden [2015a] has established an early conceptual connection between creativity and ALife. First, “psychological creativity” (involving human thought and activity) is contrasted with “biological creativity”, which is described as “the ability to generate new cells, organs, organisms, or species” and includes “the (phylogenetic) creativity of evolution, the (ontogenetic) creativity of morphogenesis, and the (autopoietic) creation of individual cells”. She then re-applies one of her most well known distinctions of creative mechanisms [Boden, 2004]: “the different sorts of surprise that we experience on encountering a creative idea correspond to different mechanisms for producing novelty. Those mechanisms mark three kinds of creativity: combinational, exploratory, and transformational” [Boden, 2015a]. More specifically:
  • Combinational creativity: the creation of new concepts or artifacts by combining features of existing ones.
  • Exploratory creativity: traversing a space of such concepts or artifacts (labelled “conceptual space”) to construct or discover new and valued instances.
  • Transformational creativity: reaching new concepts or artifacts that only become accessible by altering the rules defining the space itself “so that thoughts are now possible which previously (…) were literally inconceivable”
    [Boden, 2004, p. 6].